OpenStudy (anonymous):

Solve for y: 2 /3y + 2 = 6

OpenStudy (anonymous):

what have you done so far?

OpenStudy (anonymous):

subtract -2 from 2 and 6

OpenStudy (anonymous):

yes good that gives you \[\frac{2}{3y} = 4\] now you want y by itself what you can do is multiply y to both sides and it cancels out \[\frac{2}{3y} *y = 4*y\] \[\frac{2}{3} = 4y\] Do you know what to do know to finish it?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

well look at the last part that I type you know y by itself. What can you do to take away that 4 from the y?

OpenStudy (anonymous):

divide or subtract

OpenStudy (anonymous):

divide. By what?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Yes but remember what you do to one side you also have to do it to the other side. So you have to divide to both sides by 4.

OpenStudy (anonymous):

so I would divide 2/3 by 4?

OpenStudy (anonymous):

yes it seems a little tricky but what you have is \[\frac{2}{3} * \frac{1}{4}= \frac{2}{12}\]

OpenStudy (anonymous):

so now you have\[\frac{2}{12} =y \]

OpenStudy (anonymous):

dont forget to simplify!

OpenStudy (anonymous):

6!

OpenStudy (anonymous):

\[\frac{1}{6}= y\]

OpenStudy (anonymous):

Thanks!